Suppose 50% of the registered voters in a country are Republican. If a sample of 598 voters is selected, what is the probability that the sample proportion of Republicans will be less than 47%?

Round your answer to four decimal places.
Answer #1

Solution

Given that,

p = 0.50

1 - p = 1 - 0.50 = 0.50

n = 598

= p = 0.50

=  [p ( 1 - p ) / n] =   [(0.50 * 0.50) / 598 ] = 0.0204

P( < 0.47)

= P[( - ) / < (0.47 - 0.50) / 0.0204]

= P(z < -1.47)

Using z table,

= 0.0708